# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 4.3.1 Hyperparameter Transfer
The prohibitively high computational cost of training LLMs renders traditional hyperparameter search infeasible. Consequently, a critical open question is how to reliably transfer optimal hyperparameters (e.g., learning rate, initialization) found on small-scale proxy models to large-scale target models. yang2021tuning provide a foundational breakthrough in this domain, which utilizes the Maximal Update Parametrization ($\mu P$) to theoretically guarantee that training dynamics remain stable as model width increases, thereby enabling zero-shot hyperparameter transfer. To verify the practical limits of this theory, lingle2025empiricalstudymuplearning conducts extensive experiments, confirming the efficacy of $\mu$-Transfer while identifying crucial architectural sensitivities. Moving beyond width-based transfer, researchers have sought to establish more comprehensive laws governing hyperparameter scaling with respect to both model size and data volume. li2025predictable introduce the “Step Law,” a convex optimization framework that derives precise power-law relationships for optimal learning rates and batch sizes dependent on parameter count ($N$) and dataset size ($D$). Complementing this, filatov2025optimal identify a “norm transfer” phenomenon, proposing that the operator norm of the output layer serves as the single invariant controlling the joint optimal scaling of model and data. Finally, addressing the optimization mechanism itself, kim2025stochastic reframe the search problem, proposing a stochastic bi-level optimization algorithm that leverages Langevin dynamics to efficiently handle the uncertainty and non-convexity inherent in hyperparameter landscapes.
@@ 4.3.2 The Evolution of Optimization Algorithms
In this subsection, we focus on recent theoretical advances in optimization methods for training LLMs. In NLP tasks, Transformer-based models commonly use the Adam optimizer and its variants [vaswani2017attention, radford2019language, brown2020language]. From the perspective of optimizer development, Adam combines both first-order and second-order information [kingma2014adam]. Its first-order update comes from the momentum technique, which can be viewed as an exponential moving average of gradients [polyak1992acceleration, ruppert1988efficient]. Its second-order update is inspired by Adagrad [duchi2011adaptive] and RMSProp [hinton2012rmsprop], both of which are essentially variants of SGD. Although the original Adam paper provides a convergence proof, reddi2019convergence present a counterexample showing that Adam can fail to converge. Let $\beta_{1}$ and $\beta_{2}$ denote the hyperparameters for the first- and second-moment updates, reddi2019convergence show that when $\beta_{1}<\sqrt{\beta_{2}}$, one can construct a problem for which Adam diverges. This triggers a large body of work proposing Adam variants with guaranteed convergence. However, in practical NLP applications Adam performs very well, creating a gap between theory and practice [radford2019language, brown2020language]. zhang2022adam attempt to bridge this gap through a more refined analysis. Specifically, they show that Adam converges without any modification, as long as $\beta_{1}$ and $\beta_{2}$ are set appropriately. If $\beta_{2}$ is chosen too small, Adam will diverge. A key insight from this theory and subsequent analyses is that when the batch size is small, $\beta_{2}$ should be set to a larger value.
@@
In adversarial neural networks and reinforcement learning, researchers often use Adam instead of SGD because Adam usually shows faster convergence in practice. However, there is no definitive theoretical result proving that Adam is better than SGD. Many theoretical studies have tried to analyze Adam and SGD from different perspectives. zhang2020adaptive show, both experimentally and theoretically, that the gradient noise in Transformer-based NLP training is heavy-tailed, and such heavy-tailed noise explains why SGD performs worse than Adam. Different from stochastic gradient noise, kunstner2024heavy point out that class imbalance, which is common in language tasks, also creates heavy-tailed behavior, and this is another reason why SGD converges more slowly than Adam. wang2024convergence analyze the limitations of the uniform smoothness assumption in studying Adam’s convergence speed and introduce a non-uniform smoothness assumption. Based on this new assumption, they prove when Adam can converge faster than SGD. zhang2024transformers further observe that different parameter blocks in Transformers have heterogeneous Hessian structures. Under such block heterogeneity, SGD performs poorly because it uses the same learning rate for all parameters, while Adam’s adaptive learning rate allows it to handle heterogeneity more effectively. vasudeva2025rich study the implicit bias of Adam and SGD. Their theoretical and empirical results show that SGD exhibits a simplicity bias, which leads to weaker generalization when the data distribution changes. In contrast, Adam is more resistant to this simplicity bias and is therefore more robust under distribution shifts. ahn2024adam prove that Adam with model exponential moving average is effective for nonconvex optimization.
@@
Recent advances in LLM training have been driven largely by this observation, motivating the development of a family of non-Euclidean and matrix-aware optimizers [gupta2018shampoo]. Among these, the Muon optimizer, built on matrix orthogonalization, has emerged as a highly competitive alternative to AdamW [loshchilov2019decoupledweightdecayregularization], consistently demonstrating faster convergence and improved empirical performance in LLM training [liu2025muon]. Muon performs updates through orthogonalized momentum, a mechanism grounded in the theory of modular dualization [chen2025muon], which interprets gradients as dual-space objects that must be mapped back into the parameter’s primal space. This viewpoint provides a unified theoretical foundation for scalable training algorithms. Both Muon and related approaches such as Soap [gupta2018shampoo, vyas2024soapimprovingstabilizingshampoo] accelerate training by applying matrix-valued preconditioners—multiplying gradients by entire matrices rather than element-wise scalars.
@@
The strength of Muon is theoretically attributed to its ability to leverage the low-rank and approximately block-diagonal structure of the Hessian commonly observed in LLMs. Muon and similar spectral methods, including Spectral Descent [bernstein2024modular, bernstein2024old], also exhibit an implicit bias toward solutions maximizing margins under the spectral norm, offering potential generalization benefits. A related line of work builds on the Linear Minimization Oracle (LMO) framework [pethick2025trainingdeeplearningmodels], which includes Muon as prominent instances. PolarGrad [lau2025polargrad] further unify matrix-aware preconditioned methods by distinguishing vector-based from matrix-based preconditioning and introduce a broader class of optimizers grounded in the polar decomposition of gradient matrices, with Muon arising as a scaled nuclear-norm instance.
@@ 5.1 Fundamental Problems
These two questions concern the theoretical bounds of safety guarantees and the internal mechanisms of capability elicitation, form the bedrock of AI Alignment theory. We conclude the landscape of current theoretical consideration in fig. 5. In the following section, we review the core theories and methods the community has developed to address these profound challenges.
@@ 5.2.1 Foundations for AI Alignment
A primary line of theoretical inquiry focuses on establishing the hard limits of safety and alignment. Unlike empirical evaluations which can only show the presence of failures, these works seek to prove whether safe alignment is mathematically achievable. wolf2024fundamental implies that alignment methods that merely attenuate bad behaviors without completely removing them cannot theoretically guarantee safety against adversarial jailbreaking. Beyond individual model safety, falahati2025alignment analyze the interaction between model owners and the public under recursive filtering mechanisms. They prove an impossibility theorem, demonstrating that recursive curation cannot simultaneously satisfy diversity, fairness, and stability.
@@
Another dimension of theoretical analysis investigates how alignment modifies the underlying model, challenging the assumption that fine-tuning fundamentally alters the model’s knowledge or capabilities. ji2025language introduce the concept of “elasticity”, positing that aligned models possess a tendency to revert to their pre-training distribution. The authors theoretically derive that, compared to pre-training, the effects of alignment fine-tuning are disproportionately easily compromised. Complementing this view, other research examines the depth at which alignment operates. qi2025safety identify the phenomenon of “Shallow Safety Alignment”. The authors argue that current alignment methods essentially function as optimization shortcuts, altering only the generation distribution of the first few tokens to trigger refusal responses, while leaving the harmful knowledge in deeper layers intact.
@@
Theoretical works have also begun to uncover specific anomalies within the optimization objectives of alignment algorithms themselves. razin2025unintentional identify a critical failure mode termed “likelihood displacement”. The authors prove that this mechanism can inadvertently shift probability mass to semantically opposite responses, highlighting that standard preference optimization does not guarantee semantic alignment.
@@
Weak-to-Strong Generalization. Superalignment [openai_superalignment] is the critical challenge in AI safety of ensuring that superintelligent AI systemscan act in accordance with human values, intentions, and goals. The fundamental difficulty lies in developing a reliable mechanism to control or align an entity vastly more intelligent than its creators. This is a crucial, long-term research problem, with dedicated efforts from groups like OpenAI focused on creating technical solutions to prevent future superintelligence from going rogue or causing unintended harm to humanity. One of the core technical challenges within this framework is scalable oversight [bowman2022measuring], which seeks to enable relatively weak human supervisors to reliably evaluate and align AI systems that are far stronger and more complex than themselves [ji2023ai, shen2023large].
@@
In response to the challenge of superalignment, OpenAI introduced a well-designed paradigm termed weak-to-strong generalization (W2SG) [burns2024weaktostrong]. Their key finding demonstrates that when strong pre-trained language models are fine-tuned using supervision signals from weaker models, they consistently surpass the performance of their weak supervisors. Building upon this discovery, a growing body of research empirically investigates the properties of W2SG [yang2024super, goel2025great], and the potential of this paradigm on other tasks [guo2024vision, yang-etal-2024-weak] or scenarios [pawelczyk2024generalizing, zhou2025weak]. Additionally, various techniques are also developed to enhance the strong model’s performance in W2SG. Popular approaches include iterative updating [lyu2024macpo, ye2025iterative, lang2025debate], and incorporating more weak supervisors [agrawal2024ensemw2s, sang2024improving, liu2024co, cui2024bayesian]. In parallel, theoretical understanding of W2SG mainly focuses on whether it occurs, i.e., under what circumstances the strong student outperforms the weak teacher. Building on a convex fine-tuning function class assumption, several works [charikar2024quantifying, mulgund2025relating, yao2025understanding, yao2025revisiting] derive generalization bounds akin to the Pythagorean theorem, quantifying how much a strong student model can outperform its weak teacher via their misfit error:
@@
where $F^{\star}$ is the labeling function, $F_{w}$ is the weak model, and $F_{sw}$ is the weak-to-strong model fine-tuned with the weak label. The Kullback–Leibler (KL) divergence loss function measures the difference between two models over the data distribution, which is equivalent to the cross-entropy loss used in classification. xu2025on go further by employing bias-variance decompositions for the Bregman divergence, thereby overcoming the convexity assumption inherent in misfit-based analysis. This work demonstrates that W2SG is more likely to emerge when the student model approximates its posterior mean teacher rather than merely mimicking an individual teacher. From the perspective of a general definition of adversarial robustness, W2SG arises under appropriate data neighborhood conditions that enable weak supervision error correction [lang2024theoretical] or sufficient overlap between easy and hard patterns that allow weak supervision to guide the student in learning challenging features [shin2024weak]. Under Gaussian data assumptions, the theoretical foundations of W2SG are rigorously characterized through several frameworks: model and distribution shift [ildiz2025highdimensional], transfer learning [somerstep2025transfer] and intrinsic dimension [dong2025discrepancies]. Further theoretical insights are established through representation analysis [xue2025representations], feature learning [wu2024provable, oh2025from, moniri2025mechanisms] and random feature model [medvedev2025weak].
@@ 5.2.2 Reinforcement Learning for Alignment
The Role of RL. A central debate in the theoretical community concerns whether RL truly instills new reasoning capabilities or merely elicits latent abilities acquired during pre-training. Several studies suggest that RL primarily acts as a mechanism for efficiency and elicitation rather than capability expansion. yue2025doesreinforcementlearningreally systematically evaluate RLVR (RL with Verifiable Rewards) and argue that while RL improves sampling efficiency, it does not introduce fundamentally new reasoning patterns, with performance ultimately bounded by the base model’s distribution. shao2025spurious support this and find that even weak or random reward signals can significantly improve mathematical reasoning. The authors attribute this to the fact that RL activates valid reasoning modes (such as code-based reasoning) already present in the pre-trained model, rather than learning from the reward signal itself. zhao2025echo further characterize RL as an “echo chamber” that converges to a single dominant output format found in the pre-training data, effectively suppressing diversity while enabling positive transfer from simple to complex tasks. However, this view is also challenged by other findings. liu2025prorl demonstrate that with sufficient training duration and periodic policy resets, RL can indeed drive models to explore novel strategies absent in the base model, thereby expanding the reasoning boundary. From a geometric perspective, zhu2025path offer a theoretical explanation for these behaviors. The authors prove that RL updates occur in low-curvature subspaces orthogonal to the principal components updated by SFT. This suggests that RL operates in a distinct optimization regime, fine-tuning the model’s behavior without significantly altering its primary feature representations.
@@
Comparison of RL Paradigms. Researchers have sought to unify different RL algorithms under general frameworks. azar2024general theoretically decompose the performance gap into exact optimization and finite-sample regimes. They prove that RLHF is superior when the policy model is misspecified, whereas DPO [rafailov2023direct] excels when the reward model is misspecified. Efficiency and exploration remain critical challenges. pmlr-v267-zhong25b introduce a Reinforced Token Optimization (RTO) framework, proving that modeling RLHF as a token-wise MDP is significantly more sample-efficient than the traditional contextual bandit formulation. Meanwhile, xiong2024iterative address the lack of exploration in offline DPO. By formulating the problem as a reverse-KL regularized bandit, they propose iterative algorithms that significantly outperform static baselines.
@@
The Limits of RL. The efficacy of RL is fundamentally limited by the quality of the reward signal. “Reward hacking”, where the model exploits flaws in the reward model, is a persistent theoretical concern. Theoretical analyses on this phenomenon are pessimistic. gaikwad2025murphys introduce an alignment trilemma, mathematically proving that it is impossible to simultaneously achieve strong optimization pressure, high-fidelity value capture, and robust generalization. This is also quantified by gao2023scaling, the authors establish a functional relationship between the golden reward and the KL divergence. Crucially, they find that while increasing the reward model size improves robustness, increasing the policy model size does not mitigate overoptimization, and KL penalties act merely as early stopping mechanisms rather than true solutions. To address these vulnerabilities, miao2024inform propose a variational information bottleneck approach. By filtering out irrelevant information in the reward model’s representation, this method theoretically and empirically reduces the model’s reliance on spurious features. Finally, lin-etal-2024-mitigating discuss the trade-off between alignment and the retention of pre-training knowledge, proposing Heterogeneous Model Averaging (HMA) to balance these competing objectives.

### v2

@@ 4.3.1 Hyperparameter Transfer
The prohibitively high computational cost of training LLMs renders traditional hyperparameter search infeasible. Consequently, a critical open question is how to reliably transfer optimal hyperparameters (e.g., learning rate, initialization) found on small-scale proxy models to large-scale target models. Yang et al. (2021) provide a foundational breakthrough in this domain, which utilizes the Maximal Update Parametrization ($\mu P$) to theoretically guarantee that training dynamics remain stable as model width increases, thereby enabling zero-shot hyperparameter transfer. To verify the practical limits of this theory, Lingle (2025) conducts extensive experiments, confirming the efficacy of $\mu$-Transfer while identifying crucial architectural sensitivities. Moving beyond width-based transfer, researchers have sought to establish more comprehensive laws governing hyperparameter scaling with respect to both model size and data volume. Li et al. (2025c) introduce the “Step Law,” a convex optimization framework that derives precise power-law relationships for optimal learning rates and batch sizes dependent on parameter count ($N$) and dataset size ($D$). Complementing this, Filatov et al. (2025) identify a “norm transfer” phenomenon, proposing that the operator norm of the output layer serves as the single invariant controlling the joint optimal scaling of model and data. Finally, addressing the optimization mechanism itself, Kim and Hospedales (2025) reframe the search problem, proposing a stochastic bi-level optimization algorithm that leverages Langevin dynamics to efficiently handle the uncertainty and non-convexity inherent in hyperparameter landscapes.
@@ 4.3.2 The Evolution of Optimization Algorithms
In this subsection, we focus on recent theoretical advances in optimization methods for training LLMs. In NLP tasks, Transformer-based models commonly use the Adam optimizer and its variants (Vaswani et al., 2017b; Radford et al., 2019; Brown et al., 2020). From the perspective of optimizer development, Adam combines both first-order and second-order information (Kingma and Ba, 2014). Its first-order update comes from the momentum technique, which can be viewed as an exponential moving average of gradients (Polyak and Juditsky, 1992; Ruppert, 1988). Its second-order update is inspired by Adagrad (Duchi et al., 2011) and RMSProp (Hinton, 2012), both of which are essentially variants of SGD. Although the original Adam paper provides a convergence proof, Reddi et al. (2019) present a counterexample showing that Adam can fail to converge. Let $\beta_{1}$ and $\beta_{2}$ denote the hyperparameters for the first- and second-moment updates, Reddi et al. (2019) show that when $\beta_{1}<\sqrt{\beta_{2}}$, one can construct a problem for which Adam diverges. This triggers a large body of work proposing Adam variants with guaranteed convergence. However, in practical NLP applications Adam performs very well, creating a gap between theory and practice (Radford et al., 2019; Brown et al., 2020). Zhang et al. (2022) attempt to bridge this gap through a more refined analysis. Specifically, they show that Adam converges without any modification, as long as $\beta_{1}$ and $\beta_{2}$ are set appropriately. If $\beta_{2}$ is chosen too small, Adam will diverge. A key insight from this theory and subsequent analyses is that when the batch size is small, $\beta_{2}$ should be set to a larger value.
@@
In adversarial neural networks and reinforcement learning, researchers often use Adam instead of SGD because Adam usually shows faster convergence in practice. However, there is no definitive theoretical result proving that Adam is better than SGD. Many theoretical studies have tried to analyze Adam and SGD from different perspectives. Zhang et al. (2020) show, both experimentally and theoretically, that the gradient noise in Transformer-based NLP training is heavy-tailed, and such heavy-tailed noise explains why SGD performs worse than Adam. Different from stochastic gradient noise, Kunstner et al. (2024) point out that class imbalance, which is common in language tasks, also creates heavy-tailed behavior, and this is another reason why SGD converges more slowly than Adam. Wang et al. (2024a) analyze the limitations of the uniform smoothness assumption in studying Adam’s convergence speed and introduce a non-uniform smoothness assumption. Based on this new assumption, they prove when Adam can converge faster than SGD. Zhang et al. (2024c) further observe that different parameter blocks in Transformers have heterogeneous Hessian structures. Under such block heterogeneity, SGD performs poorly because it uses the same learning rate for all parameters, while Adam’s adaptive learning rate allows it to handle heterogeneity more effectively. Vasudeva et al. (2025) study the implicit bias of Adam and SGD. Their theoretical and empirical results show that SGD exhibits a simplicity bias, which leads to weaker generalization when the data distribution changes. In contrast, Adam is more resistant to this simplicity bias and is therefore more robust under distribution shifts. Ahn and Cutkosky (2024) prove that Adam with model exponential moving average is effective for nonconvex optimization.
@@
Recent advances in LLM training have been driven largely by this observation, motivating the development of a family of non-Euclidean and matrix-aware optimizers (Gupta et al., 2018). Among these, the Muon optimizer, built on matrix orthogonalization, has emerged as a highly competitive alternative to AdamW (Loshchilov and Hutter, 2019), consistently demonstrating faster convergence and improved empirical performance in LLM training (Liu et al., 2025c). Muon performs updates through orthogonalized momentum, a mechanism grounded in the theory of modular dualization (Chen et al., 2025a), which interprets gradients as dual-space objects that must be mapped back into the parameter’s primal space. This viewpoint provides a unified theoretical foundation for scalable training algorithms. Both Muon and related approaches such as Soap (Gupta et al., 2018; Vyas et al., 2024) accelerate training by applying matrix-valued preconditioners—multiplying gradients by entire matrices rather than element-wise scalars.
@@
The strength of Muon is theoretically attributed to its ability to leverage the low-rank and approximately block-diagonal structure of the Hessian commonly observed in LLMs. Muon and similar spectral methods, including Spectral Descent (Bernstein and Newhouse, 2025; 2024), also exhibit an implicit bias toward solutions maximizing margins under the spectral norm, offering potential generalization benefits. A related line of work builds on the Linear Minimization Oracle (LMO) framework (Pethick et al., 2025), which includes Muon as prominent instances. PolarGrad (Lau et al., 2025) further unify matrix-aware preconditioned methods by distinguishing vector-based from matrix-based preconditioning and introduce a broader class of optimizers grounded in the polar decomposition of gradient matrices, with Muon arising as a scaled nuclear-norm instance.
@@ 5.1 Fundamental Problems
These two questions concern the theoretical bounds of safety guarantees and the internal mechanisms of capability elicitation, form the bedrock of AI Alignment theory. We conclude the landscape of current theoretical consideration in fig.˜5. In the following section, we review the core theories and methods the community has developed to address these profound challenges.
@@ 5.2.1 Foundations for AI Alignment
A primary line of theoretical inquiry focuses on establishing the hard limits of safety and alignment. Unlike empirical evaluations which can only show the presence of failures, these works seek to prove whether safe alignment is mathematically achievable. Wolf et al. (2024) imply that alignment methods that merely attenuate bad behaviors without completely removing them cannot theoretically guarantee safety against adversarial jailbreaking. Beyond individual model safety, Falahati et al. (2025) analyze the interaction between model owners and the public under recursive filtering mechanisms. They prove an impossibility theorem, demonstrating that recursive curation cannot simultaneously satisfy diversity, fairness, and stability.
@@
Another dimension of theoretical analysis investigates how alignment modifies the underlying model, challenging the assumption that fine-tuning fundamentally alters the model’s knowledge or capabilities. Ji et al. (2025) introduce the concept of “elasticity”, positing that aligned models possess a tendency to revert to their pre-training distribution. The authors theoretically derive that, compared to pre-training, the effects of alignment fine-tuning are disproportionately easily compromised. Complementing this view, other research examines the depth at which alignment operates. Qi et al. (2025) identify the phenomenon of “Shallow Safety Alignment”. The authors argue that current alignment methods essentially function as optimization shortcuts, altering only the generation distribution of the first few tokens to trigger refusal responses, while leaving the harmful knowledge in deeper layers intact.
@@
Theoretical works have also begun to uncover specific anomalies within the optimization objectives of alignment algorithms themselves. Razin et al. (2025) identify a critical failure mode termed “likelihood displacement”. The authors prove that this mechanism can inadvertently shift probability mass to semantically opposite responses, highlighting that standard preference optimization does not guarantee semantic alignment.
@@
Weak-to-Strong Generalization. Superalignment (OpenAI, 2023) is the critical challenge in AI safety of ensuring that superintelligent AI systemscan act in accordance with human values, intentions, and goals. The fundamental difficulty lies in developing a reliable mechanism to control or align an entity vastly more intelligent than its creators. This is a crucial, long-term research problem, with dedicated efforts from groups like OpenAI focused on creating technical solutions to prevent future superintelligence from going rogue or causing unintended harm to humanity. One of the core technical challenges within this framework is scalable oversight (Bowman et al., 2022), which seeks to enable relatively weak human supervisors to reliably evaluate and align AI systems that are far stronger and more complex than themselves (Ji et al., 2023; Shen et al., 2023).
@@
In response to the challenge of superalignment, OpenAI introduced a well-designed paradigm termed weak-to-strong generalization (W2SG) (Burns et al., 2024). Their key finding demonstrates that when strong pre-trained language models are fine-tuned using supervision signals from weaker models, they consistently surpass the performance of their weak supervisors. Building upon this discovery, a growing body of research empirically investigates the properties of W2SG (Yang et al., 2025a; Goel et al., 2025), and the potential of this paradigm on other tasks (Guo et al., 2024; Yang et al., 2024d) or scenarios (Pawelczyk et al., 2024; Zhou et al., 2025b). Additionally, various techniques are also developed to enhance the strong model’s performance in W2SG. Popular approaches include iterative updating (Lyu et al., 2025; Ye et al., 2025; Lang et al., 2025), and incorporating more weak supervisors (Agrawal et al., 2024; Sang et al., 2024; Liu and Alahi, 2024; Cui et al., 2025). In parallel, theoretical understanding of W2SG mainly focuses on whether it occurs, i.e., under what circumstances the strong student outperforms the weak teacher. Building on a convex fine-tuning function class assumption, several works (Charikar et al., 2024; Mulgund and Pabbaraju, 2025; Yao et al., 2025b; a) derive generalization bounds akin to the Pythagorean theorem, quantifying how much a strong student model can outperform its weak teacher via their misfit error:
@@
where $F^{\star}$ is the labeling function, $F_{w}$ is the weak model, and $F_{sw}$ is the weak-to-strong model fine-tuned with the weak label. The Kullback–Leibler (KL) divergence loss function measures the difference between two models over the data distribution, which is equivalent to the cross-entropy loss used in classification. Xu et al. (2025a) go further by employing bias-variance decompositions for the Bregman divergence, thereby overcoming the convexity assumption inherent in misfit-based analysis. This work demonstrates that W2SG is more likely to emerge when the student model approximates its posterior mean teacher rather than merely mimicking an individual teacher. From the perspective of a general definition of adversarial robustness, W2SG arises under appropriate data neighborhood conditions that enable weak supervision error correction (Lang et al., 2024) or sufficient overlap between easy and hard patterns that allow weak supervision to guide the student in learning challenging features Shin et al. (2025). Under Gaussian data assumptions, the theoretical foundations of W2SG are rigorously characterized through several frameworks: model and distribution shift (Ildiz et al., 2025), transfer learning (Somerstep et al., 2025) and intrinsic dimension (Dong et al., 2025). Further theoretical insights are established through representation analysis (Xue et al., 2025), feature learning (Wu and Sahai, 2025; Oh et al., 2025; Moniri and Hassani, 2025) and random feature model (Medvedev et al., 2025).
@@ 5.2.2 Reinforcement Learning for Alignment
The Role of RL. A central debate in the theoretical community concerns whether RL truly instills new reasoning capabilities or merely elicits latent abilities acquired during pre-training. Several studies suggest that RL primarily acts as a mechanism for efficiency and elicitation rather than capability expansion. Yue et al. (2025) systematically evaluate RLVR (RL with Verifiable Rewards) and argue that while RL improves sampling efficiency, it does not introduce fundamentally new reasoning patterns, with performance ultimately bounded by the base model’s distribution. Shao et al. (2025) support this and find that even weak or random reward signals can significantly improve mathematical reasoning. The authors attribute this to the fact that RL activates valid reasoning modes (such as code-based reasoning) already present in the pre-trained model, rather than learning from the reward signal itself. Zhao et al. (2025b) further characterize RL as an “echo chamber” that converges to a single dominant output format found in the pre-training data, effectively suppressing diversity while enabling positive transfer from simple to complex tasks. However, this view is also challenged by other findings. Liu et al. (2025d) demonstrate that with sufficient training duration and periodic policy resets, RL can indeed drive models to explore novel strategies absent in the base model, thereby expanding the reasoning boundary. From a geometric perspective, Zhu et al. (2025c) offer a theoretical explanation for these behaviors. The authors prove that RL updates occur in low-curvature subspaces orthogonal to the principal components updated by SFT. This suggests that RL operates in a distinct optimization regime, fine-tuning the model’s behavior without significantly altering its primary feature representations. More recently, Li et al. (2026) formalize RLHF through the framework of algorithmic stability and build the generalization theory under the linear reward model.
@@
Comparison of RL Paradigms. Researchers have sought to unify different RL algorithms under general frameworks. Azar et al. (2024) theoretically decompose the performance gap into exact optimization and finite-sample regimes. They prove that RLHF is superior when the policy model is misspecified, whereas DPO (Rafailov et al., 2023) excels when the reward model is misspecified. Efficiency and exploration remain critical challenges. Zhong et al. (2025a) introduce a Reinforced Token Optimization (RTO) framework, proving that modeling RLHF as a token-wise MDP is significantly more sample-efficient than the traditional contextual bandit formulation. Meanwhile, Xiong et al. (2024) address the lack of exploration in offline DPO. By formulating the problem as a reverse-KL regularized bandit, they propose iterative algorithms that significantly outperform static baselines.
@@
The Limits of RL. The efficacy of RL is fundamentally limited by the quality of the reward signal. “Reward hacking”, where the model exploits flaws in the reward model, is a persistent theoretical concern. Theoretical analyses on this phenomenon are pessimistic. Gaikwad (2025) introduce an alignment trilemma, mathematically proving that it is impossible to simultaneously achieve strong optimization pressure, high-fidelity value capture, and robust generalization. This is also quantified by Gao et al. (2023), the authors establish a functional relationship between the golden reward and the KL divergence. Crucially, they find that while increasing the reward model size improves robustness, increasing the policy model size does not mitigate overoptimization, and KL penalties act merely as early stopping mechanisms rather than true solutions. To address these vulnerabilities, Miao et al. (2024) propose a variational information bottleneck approach. By filtering out irrelevant information in the reward model’s representation, this method theoretically and empirically reduces the model’s reliance on spurious features. Lin et al. (2024) further discuss the trade-off between alignment and the retention of pre-training knowledge, proposing Heterogeneous Model Averaging (HMA) to balance these competing objectives. More recently, Ouyang et al. (2025) analyze the issue of reward fairness from a resource allocation perspective, treating rewards as resources to be allocated while considering the trade-off between utility and fairness in their distribution.